OpenStudy (testing123guest):

Divide (5.6 x 10^15) by (6.4 x 10^2). Express your answer in scientific notation. A) .875 x 10^17 B) .875 x 10^13 C) 8.75 x 10^12 D) 8.75 x 10^13

OpenStudy (testing123guest):

@Anna1448 @Kitten_is_back @Luigi0210 @Nnesha @undeadknight26 @welshfella

OpenStudy (welshfella):

ok first calculate 5.6 / 6.4

OpenStudy (testing123guest):

0.875

OpenStudy (welshfella):

right now deal with the exponent part 10^15 / 10^2 = ?

OpenStudy (welshfella):

- subtract the exponents

OpenStudy (welshfella):

yes - but you need it in scientific notation

OpenStudy (testing123guest):

is it a?

OpenStudy (welshfella):

i think you are guessing try to work out the following 10^15 / 10^2 = 10^(15-2) = ?

OpenStudy (welshfella):

law of exponents x^a / x^b = x^(a-b)

OpenStudy (testing123guest):

sorry my computer crashed i am back now is the answer 10000000000000

OpenStudy (testing123guest):

@welshfella